Caffeine (molar mass: 194.2 g/mol) is a stimulant found in coffee, tea, many soft drinks and
chocolate. Elemental analysis of caffeine shows 49.47 mass % C, 5.19 mass % H, 28.86 mass % N
and 16.48 mass % O. What is the molecular formula of caffeine?
1
Expert's answer
2021-10-10T10:18:15-0400

Solution: (1) calculate the empirical formula, (2) compare "EFW" to molecular weight, (3) multiply empirical formula by proper scaling factor.

1) Calculate the empirical formula:

carbon: 49.98 g ÷ 12.011 g/mol = 4.16

hydrogen: 5.19 g ÷ 1.008 g/mol = 5.15

nitrogen: 28.85 g ÷ 14.007 g/mol = 2.06

oxygen: 16.48 g ÷ 15.999 g/mol = 1.03

carbon: 4.16 ÷ 1.03 = 4.04 = 4

hydrogen: 5.15 ÷ 1.03 = 5

nitrogen: 2.06 ÷ 1.03 = 2

oxygen: 1.03 ÷ 1.03 = 1

2) Empirical formula is C4H5N2O. The "empirical formula weight" is about 97.1, which gives a scaling factor of two.

3) The molecular formula is:

C4H5N2O times 2 = C8H10N4O2 <--- that's the molecular formula